An ultrasound unit is being used to measure a patient's heartbeat by combining the emitted 2.0 MHz signal with the sound waves reflected from the moving tissue of one point on the heart. The beat frequency between the two signals has a maximum value of 470 Hz . What is the maximum speed of the heart tissue? The speed of ultrasound waves in human tissue is 1540 m/s.
Answer #1

GIven is:-

Velocity of the source is = zero (because the source is stationary)

Frequency of emitted wave f = 2 Mhz or  2 \times 10^6Hz

and  \Delta f = 470 Hz

here Vobserver = the velocity of the tissue of a point on the heart

speed of the ultrasound waves in human tissue C=1540 m/s

Thus by using equation from Doppler effect we get,

\frac{\Delta f}{f} = \frac{2 \times v}{c}

\frac{470}{ 2 \times 10^6} = \frac{2 \times v}{1540}

v = 0.18m/s or 18cm /s This is the maximum speed of the heart tissue.